OpenStudy (anonymous):

Simplify (n^3-n)(n-2)!

OpenStudy (anonymous):

amorfide, it's (n-2)! (factorial)

OpenStudy (amorfide):

well this is embarrassing :P

OpenStudy (asnaseer):

\[\begin{align} (n^3-n)(n-2)!&=n(n^2-1)(n-2)!\\ &=n(n+1)(n-1)(n-2)!\\ &=(n+1)n(n-1)(n-2)!\\ &=(n+1)! \end{align}\]

OpenStudy (anonymous):

thanks so much asnaseer!

OpenStudy (asnaseer):

yw